OpenStudy (anonymous):

What is the solution to the rational equation (5/x^2+3x+2)-(1/x+2)=(1/3x+3)

hero (hero):

Hint: Make sure the denominators are factored: \(\large\frac{5}{(x+2)(x+1)} - \frac{1}{x+2} = \frac{1}{3(x+1)}\) Then multiply each fraction by (x+1) to get \(\large\frac{5}{(x+2)} - \frac{x+1}{x+2} = \frac{1}{3}\)

hero (hero):

Then combine fractions on the left side and cross multiply afterwards

OpenStudy (anonymous):

answers are: x = 7 x = 5/2 x = 5 x = 4

hero (hero):

Are you going to attempt to solve it this time? I think you should at least try.

hero (hero):

This is your work, not mine.

OpenStudy (anonymous):

i got -x^2+8x+12

OpenStudy (anonymous):

by cross multiplying .

hero (hero):

Did you combine fractions on the left side before cross multiply?

hero (hero):

You keep getting quadratic equations when there is no need for such.

hero (hero):

I can't imagine what it is that you are doing. Do you know how to combine two fractions?